OpenStudy (anonymous):

True or false: x3 + 8 can be factored over the set of real numbers.

OpenStudy (anonymous):

I think its true but to be honest I not sure.

OpenStudy (anonymous):

ahh... Thank you for trying.... I'm taking an online exam so I need to find someone who is sure.

OpenStudy (anonymous):

x^3 + 8 can be factored via sum of cubes... a^3 + b^3 = (a+b)(a^2 - ab + b^2) x^3 + 2^3 =

OpenStudy (anonymous):

Sooooo... true? or false? lol

OpenStudy (anonymous):

it can be factored and it does have a real solution.

OpenStudy (anonymous):

Ok. so true..............